OpenStudy (anonymous):

4x squared - x = 3 I got down to 4 ( x - 1\2) squared = 3 Now would I do 4 ( x - 1\2) squared = 3 - 1\2 ?

OpenStudy (anonymous):

and I'm completing the square

OpenStudy (anonymous):

\[4x ^{2}-x=3\]

OpenStudy (anonymous):

\[x(4x-1)=3\]

OpenStudy (anonymous):

\[x(4x-1)-3=0\]

OpenStudy (anonymous):

try from there i dont get what you had up there

OpenStudy (anonymous):

someone else was helping me solve this problem and I was told to bring out the 4 and then put x squared - x in parentheses which equaled 3 then I did 4 ( x ^2 - x + 1\4 - 1\4) = 3 4 ( x - 1\2) squared = 3 - 1\2, which turned into 4 (x + 1\2) squared = 5\2 and that's all that I had done

OpenStudy (anonymous):

no that is not true what i did is true can you solve the quadratic equation i got

OpenStudy (anonymous):

\[4x ^{2}-x-3=0\]

OpenStudy (anonymous):

and then would you use the quadratic formula?

OpenStudy (anonymous):

\[(4x+3)(x-1)\]

OpenStudy (anonymous):

you can factorize it no need for quadratic formula

OpenStudy (anonymous):

so then x = 1 and -3\4

OpenStudy (anonymous):

excellent buddy

OpenStudy (anonymous):

and I'm working 2x squared + 3x = 0, which is supposed to be solved using the quadratic formula. I just did x ( 2x +3) = 0 and made x = 0 and -3\2

OpenStudy (anonymous):

so, what should I have done to get it into a quadratic formula? Would I just make c = 0?

OpenStudy (anonymous):

write this \[2x ^{2}+3x=0\]

OpenStudy (anonymous):

which I did already

OpenStudy (anonymous):

now A=2 B=3 AND C=0

OpenStudy (anonymous):

ok, did that

OpenStudy (anonymous):

and the final answer I have is x = -3 +- the square root of 1 all over 4

OpenStudy (anonymous):

but then I checked the answer in the back of the book and the answer is 0 and -3\2, so I did something wrong

OpenStudy (anonymous):

LET ME GIVE YOU THE QF\[X=\frac{ -B+or-\sqrt{B ^{2}-4AC} }{ 2A}\]

OpenStudy (anonymous):

PUT IN YOUR VALUES and you can have BABIEES

OpenStudy (anonymous):

then that would be x = -3 + or - sqrt(-3)^2 - 4 (2)(0) all over 2 (2)